OpenStudy (anonymous):

If f(x) = 2x2 − 30, find f(4).

OpenStudy (saifoo.khan):

\[ f(x) = 2x^2 − 30 \] f(4) = 2(4)^2 − 30 \[ f(4) = 32 − 30 \] f(4) = 2

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

saifoo.khan wat happen i thought you were goin to help me